OpenStudy (anonymous):

Evaluate: Integral of (sin^2 x cos ^3 x) dx. The answer is (1/3)sin^3 x - (1/5) sin^5 x + C Need help with how to get to the answer.

OpenStudy (anonymous):

i think you can rewrite \(\cos^3(x)\) as \((1-\sin^2(x))\cos(x)\)

OpenStudy (anonymous):

then you get \[\int \sin^2(x)(1-\sin^2(x))\cos(x)dx\] and u - sub via \(u=\sin(x), du=\cos(x)dx\)

OpenStudy (anonymous):

I'll try it right now.

OpenStudy (anonymous):

ok, should work

OpenStudy (anonymous):

yep, it does. Thank you!